# Blockchain-empowered Federated Learning: Benefits, Challenges, and Solutions

### v1

@@ 4.1.2 Reinforcement Learning
Reinforcement Learning (RL) is a machine learning algorithm that enables an agent to interact with the environment, learn from its experiences, and take action accordingly. The ultimate objective is to maximize the cumulative reward obtained by the agent over time. The traditional optimization methods are ineffective in the BC-FL system because of the system’s complexity, a large number of participants, and their limited computing and communication resources. To address these challenges and achieve better results, RL can be utilized to optimize resource allocation and schedule the resources of each client based on signals received from them. This can potentially reduce system delays and lead to improved performance.
@@ 4.3 Storage Challenges and Solutions
In centralized FL systems, each client may need to receive the latest version of the global model from the parameter server before commencing local model training in each round. Subsequently, the client trains its local model based on local datasets and submits the updated local model to the parameter server for global model aggregation. Typically, all local models and the global model are stored on the parameter server during training process or deleted after use. However, the BC-FL systems require the preservation of various training-related model information within the blockchain. All computing nodes of the blockchain are used to store the entire blockchain ledger, including all versions of global and local model information, which unavoidably introduces storage redundancy to the entire system.
Most blockchain platforms currently impose strict limits on transactions or block sizes. For example, Bitcoin has a block size limit of 1 MB [124]. Although Ethereum does not have a theoretical limit on block size, its transactions are restricted due to the limited amount of gas available per block. Therefore, the storage challenge in BC-FL systems can be summarized in the following two aspects.
  • 1)
    Constrained storage capacity: the limited block size makes it difficult to store complex model parameter information.
  • 2)
    Redundant storage demands: a large amount of local model and global version information is stored in the blockchain, which brings unnecessary information redundancy and terrible storage challenges to the entire BC-FL system.
The storage challenges and related solutions in the BC-FL systems are illustrated in Fig. 11.
@@
As depicted in Fig. 11, the current landscape presents two prevailing strategies to tackle the storage challenges in BC-FL systems. The first approach entails chunking the FL models into distinct segments, which are then stored on the blockchain with constrained block size [125]. This methodology necessitates prior negotiation of a serialization plan among nodes. Subsequently, each split model’s size is logged as supplementary information within the transaction block. Gradual storage of the models on the BC-FL system is accomplished through the initiation of transactions. However, it is our assertion that such techniques possess restricted applicability and are suitable solely for systems characterized by a small number of supernodes, each endowed with robust storage capabilities capable of managing storage redundancy.
The second solution involves utilizing distributed storage technology to house the model, while retaining only the acquisition method on the blockchain [50, 67, 109, 126, 127]. For example, the InterPlanetary File System (IPFS) employs content addressing for file storage and retrieval, allowing users to access files using the hash value associated with the file [40]. In this methodology, solely the hash of the respective model finds its place on the blockchain. Additionally, Xu et al. incorporated a model producer within the system to provide download links to other nodes [73]. The blockchain then retains solely the model hashes and their corresponding download links, as part of this innovative approach. These approaches strive to address the intricate interplay between blockchain and FL requirements, paving the way for more efficient and effective storage management within BC-FL systems.

### v2

@@ 4.1.2 Reinforcement Learning
RL is a machine learning algorithm that enables an agent to interact with the environment, learn from its experiences, and take action accordingly. The ultimate objective is to maximize the cumulative reward obtained by the agent over time. The traditional optimization methods are ineffective in the BC-FL system because of the system’s complexity, a large number of participants, and their limited computing and communication resources. To address these challenges and achieve better results, RL can be utilized to optimize resource allocation and schedule the resources of each client based on signals received from them. This can potentially reduce system delays and lead to improved performance.
@@ 4.3 Storage Challenges and Solutions
The storage requirements for blockchain systems are inherently cumbersome, as each full node is required to maintain a complete backup of the entire system. This leads to a linear increase in the total storage size with the number of full nodes. In FL, clients transmit their local models to a central server and download the global model. The server is responsible for storing both the local and global models and various FL-related data and parameters, thereby becoming the node with the highest storage demand in the FL system. When enhancing the FL system with blockchain, the BC-FL system must inevitably store diverse information on the blockchain, resulting in significant storage overhead. Furthermore, most blockchain platforms currently impose limitations on transaction or block size. For instance, Bitcoin has a block size limit of 1MB, and while Ethereum does not have a theoretical block size limit, its gas limit effectively restricts the size of transactions [143]. If the BC-FL system requires direct storage of large volumes of data within blocks, such as model parameters, this could surpass the blockchain system’s storage capabilities.
As illustrated in Fig. 11, the storage challenges of the BC-FL system are primarily twofold:
Constrained storage capacity: the limited block size makes storing some data that takes up storage space difficult.
Redundant storage demands: a large amount of training-related data is stored in the blockchain, which brings unnecessary information redundancy and terrible storage challenges to the entire BC-FL system.
@@
As depicted in Fig. 11, the current landscape presents two prevailing strategies to tackle the storage challenges in BC-FL systems. The first approach entails chunking the FL models or data into distinct segments, which are then stored on the blockchain with constrained block size [144]. This methodology necessitates prior negotiation of a serialization plan among nodes. Subsequently, each split data’s size is logged as supplementary information within the transaction block. Gradual storage of the data on the BC-FL system is accomplished through the initiation of transactions. However, we assert that such techniques possess restricted applicability and are suitable solely for systems characterized by a few supernodes, each endowed with robust storage capabilities capable of managing storage redundancy.
The second solution involves utilizing distributed storage technology to house the model, while retaining only the acquisition method on the blockchain [60, 83, 129, 145, 146]. For example, the InterPlanetary File System (IPFS) employs content addressing for file storage and retrieval, allowing users to access files using the hash value associated with the file [50]. In this methodology, solely the hash of the respective model finds its place on the blockchain. Additionally, Xu et al. incorporated a model producer within the system to provide download links to other nodes [91]. The blockchain then retains the model hashes and corresponding download links solely as part of this innovative approach. These approaches address the intricate interplay between blockchain and FL requirements, paving the way for more efficient and effective storage management within BC-FL systems.